# Human as AI Mentor: Enhanced Human-in-the-loop Reinforcement Learning for Safe and Efficient Autonomous Driving

### 10 Jan 2024 v2

@@ 1.1 IL and RL Methods for Driving Policy Learning
Driving policy learning is the main foucs of this paper. Within the domain of learning-based driving policy learning, there are two main branches: Imitation Learning (IL) and Reinforcement Learning (RL) (Aradi, 2020). IL aims to emulate human driving behavior by replicating demonstrated actions in specific situations, such as behavior cloning (BC) (Le Mero et al., 2022), inverse RL (IRL) (Huang et al., 2023b), and generative adversarial imitation learning (GAIL) (Ho and Ermon, 2016). During the demonstration phase, the novice agent avoids interacting with potentially risky environments, thereby ensuring training safety. However, two primary inherent problems are exposed: Firstly, distributional shift (Codevilla et al., 2019), where accumulated errors over time cause a deviation from the training distribution, leading to control failure. Secondly, the limitations of asymptotic performance (Silver et al., 2018), which are constrained by the proficiency of the demonstration source and unlikely to outperform it. In this context, conservative Q-learning (CQL) (Kumar et al., 2020) draws wide attention as it addresses the impact of out-of-distribution (OOD) actions in distributional shift problem. Conversely, RL employs a self-optimizing approach with the potential to address these imitation-related pitfalls (Aradi, 2020). RL-based methods, including proximal policy optimization (PPO) and soft actor-critic (SAC), have demonstrated successful applications across numerous challenging driving scenarios in AVs (Kiran et al., 2021).
@@ 1.3 Contributions
Compared to traditional methods (e.g., IL, safe RL, and conventional human-in-the-loop RL), comparative experiments indicate that our method achieves superior driving safety, reduced traffic flow disturbance, heightened sampling efficiency, and significant generalizability across various traffic environments  [^1] Experimental demo videos are available at: https://www.youtube.com/playlist?list=PL-EmC8vF-RSH2j09uxZeyiVVePuA3Eud. Our contributions are four fold:
@@
The rest of this paper is structured as follows: Section 2 provides the preliminaries. Section 3 illustrate the inspiration behind HAIM and its two key mechanisms. Section 4 describes the HAIM-DRL framework in detail, including the reward-free actor-critic architecture and the value and policy networks within it. Section 5 presents the experimental setup and results, including comparative and ablation experiments. Section 6 serves as the conclusion of the paper and discusses potential future research directions in this field.
@@ 3.2 Details of the HAIM Paradigm
Drawing inspiration from the human learning process, we propose an innovative learning paradigm, i.e., HAIM, addressing limitation (a) we previously mentioned. In particular, the human mentor not only provides high-quality demonstrations and ensure the safety of the agent’s exploration but also instructs the agent to consider system level performance. To achieve this goal, we have carefully designed explicit and implicit intervention mechanisms.
@@ 3.2.1 Explicit Intervention Mechanism
The role of the explicit intervention mechanism is similar to that of an instructor in the driving school scenario. Specifically, human expert provides guidance, supervision, and direct intervention during the agent’s learning process. For this purpose, we adopt a technique termed the ‘switching function’ (Peng et al., 2022), which allows the agent to dynamically alternate between exploration and explicit intervention.
@@ 3.2.2 Implicit Intervention Mechanism
By using Eqs. 7 and 8, we can calculate the average velocity of the following HVs at the current time $t$, which reflects the current traffic flow state. Additionally, we are able to predict the possible traffic flow state that could arise following the execution of the current action $a_{\_}t^{AV}$. In the following sections, we will integrate these two indicators as components of the disturbance cost into the RL training process. This will guide AVs in making decisions that balance its own performance with the system-level performance.
@@ 4.2.2 Removing the Reward Function
In a recent survey (Booth et al., 2023), it was found that 92% of surveyed expert-level RL practitioners reported using manual trial-and-error reward designs, with 89% indicating that their designed reward were suboptimal. Knox et al. (2023) developed eight sanity checks to identify flaws in reward functions and surveyed papers on RL-based AV published in top-tier venues, finding near-universal flaws in reward design for AVs. Thus, designing a reward function that is universally applicable to all scenarios for AVs is a complex and challenging task. Recent work has demonstrated that agents can learn complex goals by interacting with humans without a fixed reward function (Li et al., 2022c; Ibarz et al., 2018; Peng et al., 2023).
@@ 4.3.3 Learning from Implicit Intervention
Eq. 2 illustrates that even minor changes in the state of the leading AV can lead to disturbances in the following HVs, disrupting the overall traffic flow balance. In the transportation community, there are many indicators used to assess the performance of the traffic system, such as acceleration, headway, average velocity, and time to collision (TTC) (Shi et al., 2023; Jiang et al., 2022; Zhu et al., 2020; Mahdinia et al., 2021). In particular, Sharma et al. (2021) proposed using a wavelet energy-based method to calculate traffic oscillation metrics (i.e., oscillation duration and amplitude) to analyze traffic flow disturbance. The main idea of their work is to identify significant changes in the velocity. Inspired by this, and building upon Eq. 3, we propose a measurement method for disturbance cost, which can quantify the potential downstream impact of AVs on the following vehicles.
@@
This exponential formulation of the disturbance cost ensures that it increases with the deviation of the actual average velocity of the following vehicles from their desired velocity, thus encouraging the AV to behave in a way that minimizes this deviation. The desired velocity calculation of the following vehicles is closely tied to the current velocity of the leading AV. Thus, disturbance cost is closely related to the current velocity of the AV. For instance, if the throttle is depressed to the same degree, a lower velocity of the AV should result in a relatively smaller cost.
@@ 4.3.4 Reducing the Human Mentor’s Cognitive Load
The learning agent faces the risk of excessive dependence on human if the frequency of human takeover is not restricted, leading to failure when evaluating the agent without human supervision (Peng et al., 2022; Li et al., 2022c; Wu et al., 2023). This risk arises because $\hat{Q}(s_{\_}t,a_{\_}t)$ represents the proxy $Q$ value of the mixed behavior policy $\pi_{\_}{mix}$, rather than the AV’s learning policy $\pi_{\_}{AV}$, owing to the presence of human takeover. In this case, there’s a likelihood that the agent may select actions that go against human wishes, such as veering off the road when approaching the boundary. This behavior compels the human to intervene and offer corrective demonstrations. In consequence, the AV’s degree of automation remains low, leading to the human mentor experiencing fatigue due to the continual need to provide demonstrations.
@@ 5.1.4 Implementation Details
As emphasized in Assumption 1, the HAIM-DRL framework requires only a single experienced human expert for implementation. In this experiment, as shown in Figs. 5 (a), a human participant with both United States and Chinese driving licenses and 7 years of driving experience was invited to supervise the learning agent’s real-time exploration with hands on the steering wheel. Upon perceiving an impending dangerous situation, the human takeover by pressing the paddle beside the wheel, thereby gaining control of the vehicle through steering and pedal manipulation. Additionally, we assume that there are $N$ HVs following the AV agent, driving according to the IDM. Too large an ‘$N$’ can complicate traffic management, whereas too small an ‘$N$’ might oversimplify real-world traffic dynamics. Similarly, a very high ‘$X$’ can overwhelm the AV with unpredictable elements, while a low ‘$X$’ may not adequately test its adaptability. In our study, the ‘$N$’ value aligns with Shi et al. (2023), and ‘$X$’ with Li et al. (2022c). The B describes the detailed hyperparameter settings used in this experiment.
@@ 5.1.5 Training Environment
We divided the generated driving scenarios into two sets: a training set and a test set. The AV agent were exclusively trained on the training set and subsequently evaluated on the reserved test set. At the outset of each episode, we randomly selected a scenario from either the training set or the test set. Following each training iteration, we deployed the learned agent into the test environment without any human supervision. Experiments are carried out using the MetaDrive/CARLA simulator, and methods are implemented with RLLib (Liang et al., 2018), an efficient distributed learning system.
@@ 5.2 Comparative Results in the MetaDrive Simulator
In the comparative analysis, each table and each figure reports the standard deviation of multiple replicated experiments, each using a different randomized seed. The data presented in Table 1 represent the average of the maximum values observed at the last checkpoint of each evaluation, following the protocol of repeating the experiments five times to ensure the robustness and reliability of the results.
@@ 5.2.1 Compared to RL and Safe RL Methods
Safety is our primary concern, and HAIM-DRL significantly reduces safety violation. It records a mere 29.84 safety violation during the entire training phase. This figure is significantly lower, by two orders of magnitude, compared to other RL methods, despite not incorporating environmental cost. We find that the RS techniques in SAC and PPO variants are less effective than the Lagrangian methods. Furthermore, PPO exhibits more violation than SAC, which might be attributable to its relatively lower sampling efficiency and slower convergence rate. Regarding generalization ability, HAIM-DRL achieves a notable test success rate of 0.85 after with a safety violation rate of just 0.76. This achievement not only showcases its potential for navigating to destinations in new and uncertain driving environments but also ensures the AV’s safety amid these environmental uncertainties. This result is pivotal as it has the potential to increase public trust in AV technologies.
@@ 5.2.4 Compared to Human-AI Copilot Method
We also conducted a comparative analysis with the recently developed human-AI copilot optimization (HACO) method(Li et al., 2022c). We used the average episodic intervention occurrence to represent the frequency of interventions. As indicated in the Fig. 9, at the start of the training, the human expert is more frequently involved in driving demonstrations to prevent the agent from entering hazardous states. Subsequently, as the agent acquire basic driving skills, it tend to select actions more agreeable to the human expert, thereby reducing takeover rate. The takeover cost and takeover rate curves show that both HACO and HAIM-DRL experience a decline in human takeover frequency, while the consistency between agent and human behavior increases. This decline in HAIM-DRL can be attributed to its intervention minimization mechanism, which reduces both the cost and frequency of human takeover.
@@
Importantly, HACO did not take into account the impact on subsequent traffic flow, leading to fluctuations in trffic flow. As shown in Fig. 9, at the 20k step, HAIM-DRL’s disturbance rate already stabilizes at around 0.02 with a disturbance cost of about 10, while HACO’s disturbance rate is still fluctuating around 0.04, with a disturbance cost of approximately 15. To further illustrate the advantages of HAIM-DRL in terms of traffic flow efficiency, we selected an episode for a detailed presentation, as shown in Fig. 10. The blue curve represents the trajectories of both HACO and HAIM-DRL, while the other curves represent IDM-controlled HVs following the AV. The difference in the horizontal axis lengths between the HACO and HAIM-DRL plots is attributed to the higher average velocity of HAIM-DRL, which enables the vehicles to reach their destination more quickly.
It is evident that the HACO-controlled AV exhibits more frequent stopping behaviors, whereas the trajectory of HAIM-DRL is smoother. In terms of velocity, HACO leads to three abrupt drops to zero, while HAIM-DRL experiences only one. This stop for HAIM-DRL occurs at a roundabout, necessitating a yield, which is reasonable. Notably, the relatively low upper-velocity limit is due to restrictions on the maximum velocity of steering wheel-controlled vehicles to 40 km/h in the MetaDrive simulator. From the Fig. 10 (c), HACO sometimes decelerate sharply, exceeding -10, causing traffic flow disturbance (indicated by the dense purple area). In contrast, HAIM-DRL, due to its implicit intervention mechanism, employs gentler braking, thus causing less disturbance to subsequent traffic flow. Overall, HAIM-DRL not only ensures safe driving for AVs but also positions them as courteous and considerate drivers within the traffic system.
@@ 5.3.1 Analysis of Explicit Intervention Strategy
We engaged a human participant to test two different takeover strategies. The first strategy involves a small number of takeover, each producing long trajectories and typically generating sparse takeover cost. The second strategy requires more frequent takeover, leading to fragmented demonstrations. As illustrated in Table 2 (a), HAIM-DRL performs best with dense human takeover signals. AVs trained with longer trajectories, in contrast to those receiving dense takeover signals, demonstrate lower success rates and experience reward in a more erratic manner.
@@ 5.4 Comparative Results in the CARLA Simulator
All three methods show a decline in test success rate in CARLA compared to their performance in the MetaDrive. This difference is primarily attributed to the variance in training sample sizes. While HAIM-DRL is trained with 30K samples in MetaDrive, its training in CARLA is limited to just 8K samples. This significant reduction in the number of training samples inevitably impacts the methods’ ability to learn and adapt to the driving environment. Besides, in CARLA, the training is confined to a single scenario, limiting the scope and diversity of the learning experience. In contrast, MetaDrive offers a virtually infinite array of non-repetitive scenarios, providing a richer and more varied training environment.
@@ 6 Conclusions and Future Work
Despite the significant advancements in autonomous driving technology, ensuring the safety of autonomous vehicles (AVs) and traffic flow efficiency in mixed traffic environments remains a complex challenge. Addressing this, our paper propose an enhanced human-in-the-loop reinforcement learning method for driving policy learning, leveraging the synergies between the transportation and robotics communities. At the core of our approach is an innovative learning paradigm inspired by the human learning process, which is termed human as AI mentor (HAIM). A defining feature of HAIM is that the human expert serves as a mentor to the AI agent, supervising, intervening, and demonstrating in its learning process. Building upon this concept, we propose a HAIM-based deep reinforcement learning (DRL) framework, named HAIM-DRL. By using HAIM-DRL, not only can we ensure the safety of the AV itself, but also optimize the traffic flow efficiency.

### 19 Feb 2024 v3

@@ 1.1 IL and RL Methods for Driving Policy Learning
Driving policy learning is the main focus of this paper. Within the domain of learning-based driving policy learning, there are two main branches: Imitation Learning (IL) and Reinforcement Learning (RL) (Aradi, 2020). IL aims to emulate human driving behavior by replicating demonstrated actions in specific situations, such as behavior cloning (BC) (Le Mero et al., 2022), inverse RL (IRL) (Huang et al., 2023b), and generative adversarial imitation learning (GAIL) (Ho and Ermon, 2016). During the demonstration phase, the novice agent avoids interacting with potentially risky environments, thereby ensuring training safety. However, two primary inherent problems are exposed: Firstly, distributional shift (Codevilla et al., 2019), where accumulated errors over time cause a deviation from the training distribution, leading to control failure. Secondly, the limitations of asymptotic performance (Silver et al., 2018), which are constrained by the proficiency of the demonstration source and unlikely to outperform it. In this context, conservative Q-learning (CQL) (Kumar et al., 2020) draws wide attention as it addresses the impact of out-of-distribution (OOD) actions in distributional shift problem. Conversely, RL employs a self-optimizing approach with the potential to address these imitation-related pitfalls (Aradi, 2020). RL-based methods, including proximal policy optimization (PPO) and soft actor-critic (SAC), have demonstrated successful applications across numerous challenging driving scenarios in AVs (Kiran et al., 2021).
@@ 1.3 Contributions
Compared to traditional methods (e.g., IL, safe RL, and conventional human-in-the-loop RL), comparative experiments indicate that our method achieves superior driving safety, reduced traffic flow disturbance, heightened sampling efficiency, and significant generalizability across various traffic environments  [^1] Experimental demo videos are available at: https://www.youtube.com/playlist?list=PL-EmC8vF-RSH2j09uxZeyiVVePuA3Eud. Our contributions are four-fold:
@@
The rest of this paper is structured as follows: Section 2 provides the preliminaries. Section 3 illustrates the inspiration behind HAIM and its two key mechanisms. Section 4 describes the HAIM-DRL framework in detail, including the reward-free actor-critic architecture and the value and policy networks within it. Section 5 presents the experimental setup and results, including comparative and ablation experiments. Section 6 serves as the conclusion of the paper and discusses potential future research directions in this field.
@@ 3.2 Details of the HAIM Paradigm
Drawing inspiration from the human learning process, we propose an innovative learning paradigm, i.e., HAIM, addressing limitation (a) we previously mentioned. In particular, the human mentor not only provides high-quality demonstrations and ensures the safety of the agent’s exploration but also instructs the agent to consider system level performance. To achieve this goal, we have carefully designed explicit and implicit intervention mechanisms.
@@ 3.2.1 Explicit Intervention Mechanism
The role of the explicit intervention mechanism is similar to that of an instructor in the driving school scenario. Specifically, human experts provide guidance, supervision, and direct intervention during the agent’s learning process. For this purpose, we adopt a technique termed the ‘switching function’ (Peng et al., 2022), which allows the agent to dynamically alternate between exploration and explicit intervention.
@@ 3.2.2 Implicit Intervention Mechanism
By using Eqs. 7 and 8, we can calculate the average velocity of the following HVs at the current time $t$, which reflects the current traffic flow state. Additionally, we are able to predict the possible traffic flow state that could arise following the execution of the current action $a_{\_}t^{AV}$. In the following sections, we will integrate these two indicators as components of the disturbance cost into the RL training process. This will guide AVs in making decisions that balance their own performance with the system-level performance.
@@ 4.2.2 Removing the Reward Function
In a recent survey (Booth et al., 2023), it was found that 92% of surveyed expert-level RL practitioners reported using manual trial-and-error reward designs, with 89% indicating that their designed reward was suboptimal. Knox et al. (2023) developed eight sanity checks to identify flaws in reward functions and surveyed papers on RL-based AV published in top-tier venues, finding near-universal flaws in reward design for AVs. Thus, designing a reward function that is universally applicable to all scenarios for AVs is a complex and challenging task. Recent work has demonstrated that agents can learn complex goals by interacting with humans without a fixed reward function (Li et al., 2022c; Ibarz et al., 2018; Peng et al., 2023).
@@ 4.3.3 Learning from Implicit Intervention
Eq. 2 illustrates that even minor changes in the state of the leading AV can lead to disturbances in the following HVs, disrupting the overall traffic flow balance. In the transportation community, there are many indicators used to assess the performance of the traffic system, such as acceleration, headway, average velocity, and time to collision (TTC) (Shi et al., 2023; Jiang et al., 2022; Zhu et al., 2020; Mahdinia et al., 2021). In particular, Sharma et al. (2021) proposed using a wavelet energy-based method to calculate traffic oscillation metrics (i.e., oscillation duration and amplitude) to analyze traffic flow disturbance. The main idea of their work is to identify significant changes in velocity. Inspired by this, and building upon Eq. 3, we propose a measurement method for disturbance cost, which can quantify the potential downstream impact of AVs on the following vehicles.
@@
This exponential formulation of the disturbance cost ensures that it increases with the deviation of the actual average velocity of the following vehicles from their desired velocity, thus encouraging the AV to behave in a way that minimizes this deviation. The desired velocity calculation of the following vehicles is closely tied to the current velocity of the leading AV. Thus, disturbance cost is closely related to the current velocity of the AV. For instance, if the throttle is depressed to the same degree, a lower velocity of the AV would result in a relatively smaller cost.
@@ 4.3.4 Reducing the Human Mentor’s Cognitive Load
The learning agent faces the risk of excessive dependence on human if the frequency of human takeover is not restricted, leading to failure when evaluating the agent without human supervision (Peng et al., 2022; Li et al., 2022c; Wu et al., 2023). This risk arises because $\hat{Q}(s_{\_}t,a_{\_}t)$ represents the proxy $Q$ value of the mixed behavior policy $\pi_{\_}{mix}$, rather than the AV’s learning policy $\pi_{\_}{AV}$, owing to the presence of human takeover. In this case, there’s a likelihood that the agent may select actions that go against human wishes, such as veering off the road when approaching the boundary. This behavior compels human to intervene and offer corrective demonstrations. In consequence, the AV’s degree of automation remains low, leading to the human mentor experiencing fatigue due to the continual need to provide demonstrations.
@@ 5.1.4 Implementation Details
As emphasized in Assumption 1, the HAIM-DRL framework requires only a single experienced human expert for implementation. In this experiment, as shown in Figs. 5 (a), a human participant with both United States and Chinese driving licenses and 7 years of driving experience was invited to supervise the learning agent’s real-time exploration with hands on the steering wheel. Upon perceiving an impending dangerous situation, the human takes over by pressing the paddle beside the wheel, thereby gaining control of the vehicle through steering and pedal manipulation. Additionally, we assume that there are $N$ HVs following the AV agent, driving according to the IDM. Too large an ‘$N$’ can complicate traffic management, whereas too small an ‘$N$’ might oversimplify real-world traffic dynamics. Similarly, a very high ‘$X$’ can overwhelm the AV with unpredictable elements, while a low ‘$X$’ may not adequately test its adaptability. In our study, the ‘$N$’ value aligns with Shi et al. (2023), and ‘$X$’ with Li et al. (2022c). The B describes the detailed hyperparameter settings used in this experiment.
@@ 5.1.5 Training Environment
We divided the generated driving scenarios into two sets: a training set and a test set. The AV agent was exclusively trained on the training set and subsequently evaluated on the reserved test set. At the outset of each episode, we randomly selected a scenario from either the training set or the test set. Following each training iteration, we deployed the learned agent into the test environment without any human supervision. Experiments are carried out using the MetaDrive/CARLA simulator, and methods are implemented with RLLib (Liang et al., 2018), an efficient distributed learning system.
@@ 5.2 Comparative Results in the MetaDrive Simulator
In the comparative analysis, each table and each figure report the standard deviation of multiple replicated experiments, each using a different randomized seed. The data presented in Table 1 represent the average of the maximum values observed at the last checkpoint of each evaluation, following the protocol of repeating the experiments five times to ensure the robustness and reliability of the results.
@@ 5.2.1 Compared to RL and Safe RL Methods
Safety is our primary concern, and HAIM-DRL significantly reduces safety violations. It records a mere 29.84 safety violation during the entire training phase. This figure is significantly lower, by two orders of magnitude, compared to other RL methods, despite not incorporating environmental cost. We find that the RS techniques in SAC and PPO variants are less effective than the Lagrangian methods. Furthermore, PPO exhibits more violation than SAC, which might be attributable to its relatively lower sampling efficiency and slower convergence rate. Regarding generalization ability, HAIM-DRL achieves a notable test success rate of 0.85 after with a safety violation rate of just 0.76. This achievement not only showcases its potential for navigating to destinations in new and uncertain driving environments but also ensures the AV’s safety amid these environmental uncertainties. This result is pivotal as it has the potential to increase public trust in AV technologies.
@@ 5.2.4 Compared to Human-AI Copilot Method
We also conducted a comparative analysis with the recently developed human-AI copilot optimization (HACO) method(Li et al., 2022c). We used the average episodic intervention occurrence to represent the frequency of interventions. As indicated in Fig. 9, at the start of the training, the human expert is more frequently involved in driving demonstrations to prevent the agent from entering hazardous states. Subsequently, as the agent acquires basic driving skills, it tends to select actions more agreeable to the human expert, thereby reducing takeover rate. The takeover cost and takeover rate curves show that both HACO and HAIM-DRL experience a decline in human takeover frequency, while the consistency between agent and human behavior increases. This decline in HAIM-DRL can be attributed to its intervention minimization mechanism, which reduces both the cost and frequency of human takeover.
@@
Importantly, HACO did not take into account the impact on subsequent traffic flow, leading to fluctuations in traffic flow. As shown in Fig. 9, at the 20k step, HAIM-DRL’s disturbance rate already stabilizes at around 0.02 with a disturbance cost of about 10, while HACO’s disturbance rate is still fluctuating around 0.04, with a disturbance cost of approximately 15. To further illustrate the advantages of HAIM-DRL in terms of traffic flow efficiency, we selected an episode for a detailed presentation, as shown in Fig. 10. The blue curve represents the trajectories of both HACO and HAIM-DRL, while the other curves represent IDM-controlled HVs following the AV. The difference in the horizontal axis lengths between the HACO and HAIM-DRL plots is attributed to the higher average velocity of HAIM-DRL, which enables the vehicles to reach their destination more quickly.
It is evident that the HACO-controlled AV exhibits more frequent stopping behaviors, whereas the trajectory of HAIM-DRL is smoother. In terms of velocity, HACO leads to three abrupt drops to zero, while HAIM-DRL experiences only one. This stop for HAIM-DRL occurs at a roundabout, necessitating a yield, which is reasonable. Notably, the relatively low upper-velocity limit is due to restrictions on the maximum velocity of steering wheel-controlled vehicles to 40 km/h in the MetaDrive simulator. From Fig. 10 (c), HACO sometimes decelerates sharply, exceeding -10, causing traffic flow disturbance (indicated by the dense purple area). In contrast, HAIM-DRL, due to its implicit intervention mechanism, employs gentler braking, thus causing less disturbance to subsequent traffic flow. Overall, HAIM-DRL not only ensures safe driving for AVs but also positions them as courteous and considerate drivers within the traffic system.
@@ 5.3.1 Analysis of Explicit Intervention Strategy
We engaged a human participant to test two different takeover strategies. The first strategy involves a small number of takeovers, each producing long trajectories and typically generating sparse takeover cost. The second strategy requires more frequent takeover, leading to fragmented demonstrations. As illustrated in Table 2 (a), HAIM-DRL performs best with dense human takeover signals. AVs trained with longer trajectories, in contrast to those receiving dense takeover signals, demonstrate lower success rates and experience reward in a more erratic manner.
@@ 5.4 Comparative Results in the CARLA Simulator
All three methods show a decline in test success rate in CARLA compared to their performance in MetaDrive. This difference is primarily attributed to the variance in training sample sizes. While HAIM-DRL is trained with 30K samples in MetaDrive, its training in CARLA is limited to just 8K samples. This significant reduction in the number of training samples inevitably impacts the methods’ ability to learn and adapt to the driving environment. Besides, in CARLA, the training is confined to a single scenario, limiting the scope and diversity of the learning experience. In contrast, MetaDrive offers a virtually infinite array of non-repetitive scenarios, providing a richer and more varied training environment.
@@ 6 Conclusions and Future Work
Despite the significant advancements in autonomous driving technology, ensuring the safety of autonomous vehicles (AVs) and traffic flow efficiency in mixed traffic environments remains a complex challenge. Addressing this, our paper proposes an enhanced human-in-the-loop reinforcement learning method for driving policy learning, leveraging the synergies between the transportation and robotics communities. At the core of our approach is an innovative learning paradigm inspired by the human learning process, which is termed human as AI mentor (HAIM). A defining feature of HAIM is that the human expert serves as a mentor to the AI agent, supervising, intervening, and demonstrating in its learning process. Building upon this concept, we propose a HAIM-based deep reinforcement learning (DRL) framework, named HAIM-DRL. By using HAIM-DRL, not only can we ensure the safety of the AV itself, but also optimize the traffic flow efficiency.